# von Neumann entropy and quantum version of thermodynamic entropy

### v1

@@ III.1 As basis-independent representation of quantum statistical entropy
One way to understand the VN entropy is to view it as the basis-independent representation of the quantum statistical entropy of a thermodynamic system in equilibrium. The VN entropy for an appropriate density matrix can give us the qBN or qGB entropy. In fact, while deriving the expression of the thermodynamic (TH) entropy of a subsystem, Landau also came up with the same expression appears in Eq. 3 as the basis-independent representation of the quantum statistical entropy [9]. We note that, if we take $\rho=\frac{1}{d}I$, where $d$ is the dimension of the system’s relevant subspace (of the total Hilbert space) and $I$ is the identity matrix, then it is easy to check that $S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=-k_{B}\ln d=S_{qBN}$. On the other hand, if we take $\rho=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$, where $H$ the system Hamiltonian and $\beta$ is the inverse temperature, then it is difficult to verify that $S_{VN}=S_{qGB}$.
@@ III.2 As entropy of thermodynamic ensemble
Sometimes we may not know in which (pure) state our quantum system is in; von Neumann introduced the concept of the density matrix ($\rho$) to describe the physics of the system [8]. The density matrix is a statistical mixture of many pure states; for example,
@@
While following the Gibbsian way, one immediately faces problems in defining entropy of the system in the state $\rho$. First of all, the decomposition in Eq. 4 is not unique and, then, the pure states {|$\{\ket{\psi_{i}}\}$ in a decomposition may not be in general distinguishable (as they are not in general orthonormal). Through thermodynamic considerations, von Neumann proposed a formula (as given in Eq. 3) to quantify entropy of a system in state $\rho$. Without going through the details of his argument, we briefly discuss here the two main features of the VN entropy. First we distinguish two different processes: (1) time evaluation of the system, and (2) measurement on the system. If $\rho(0)$ is the density matrix at time $t=0$ then the density matrix at the $t$ is given by $\rho(t)=U(t)\rho(0)U^{{\dagger}}(t)$, where $U(t)$ is the time evaluation operator. It is not difficult to see that the VN entropy is invariant under this time evaluation process, i.e., $S_{VN}(t)=S_{VN}(0)$. von Neumann showed that a measurement, in general, irreversibly changes the density matrix to another density matrix $\rho\to\rho^{\prime}$, where
@@
We may here note that the set of states {|$\{\ket{\phi_{n}}\}$ are the eigenkets of the observable for which one is performing the measurement. The subtleties of degeneracy can be found in the von Neumann’s work [8]. He showed that the entropy of a system, in general, increases (or remains constant but never decreases) due to the measurement process, i.e., $S_{VN}(\rho^{\prime})\geq S_{VN}(\rho)$. If we identify the first process (time evaluation) and the second process (measurement) with the thermodynamic reversible and irreversible processes, respectively, then these results are consistent with the thermodynamic laws. Before we discuss, in the next subsection, some of the difficulties with the VN entropy formula, we try to intuitively understand the formula below.
As in Sec. II.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero. In this case, as we see from Eq. 3, the VN entropy is also zero. When the system is not in pure state, rather in a statistical mixture of pure states (represented by $\rho$), a formula for entropy must quantify the amount of “mixedness” (or statistical impurity) in the state $\rho$, similar to how the BN entropy quantifies the amount of “uncertainty” of a system to be found in a microstate. If we now see the decomposition in Eq. 4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-k_{B}\sum_{i}p_{i}\ln p_{i}$, which is the Shannon entropy for the probability distribution $\{p_{i}\}$. Unfortunately, the decomposition of a mixed state is not unique, hence this way of defining entropy is unacceptable. One way to circumvent this problem is to consider all possible decompositions of $\rho$ and take the one for which $S_{SN}(\{p_{i}\})$ is minimum. The advantage of defining entropy this way is that the entropy is now unique for a given $\rho$. In fact it can be proved that the VN entropy of a given $\rho$ corresponds to the aforementioned minimum [15, 16]:
@@
Another way to understand the VN entropy is as follows. If we take a set of orthonormal states to be a set of distinguishable microstates, we can define an entropy from the probabilities associated with these microstates. If {|$\{\ket{\phi_{n}}\}$ is a set of orthonormal states, the entropy one can define is $S_{SN}(\{w_{i}\})=-k_{B}\sum_{n}w_{n}\ln w_{n}$, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. Unfortunately, again, this entropy is not unique as it depends on the set of orthonormal states (microstates). To understand how to avoid this problem, we first define a new density matrix $\rho^{\prime}=\sum_{n}w_{n}\ket{\phi_{n}}\bra{\phi_{n}}$ |, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. We may note that |$\ket{\phi_{n}}$ is an eigenket of $\rho^{\prime}$ with eigenvalue $w_{n}$. This density matrix ($\rho^{\prime}$) is in the same form as the one defined in Eq. 5. But for this updated density matrix after measurement ($\rho\to\rho^{\prime}$), von Neumann showed that $-\text{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\text{Tr}(\rho\ln\rho)$ [8]. Thus we see that the VN entropy corresponds to the minimum of $S_{SN}(\{w_{i}\})$, which we obtain when we choose the microstates as the eigenkets of the given density matrix $\rho$. Accordingly,
@@
III.3 Criticism against VN entropy: Are we always asking legitimate questions?
@@
Time invariance of VN entropy:
@@
The first of these objections is that the VN entropy of an isolated quantum system does not change (more specifically, increase) as the system evolves with time. As seen earlier, $S_{VN}(t)=S_{VN}(0)$ as $\rho(0)\to\rho(t)$ under the time evaluation operator $U(t)$. It is known that, for an isolated system, the TH entropy increases as the system approaches thermal equilibrium. Does not the VN entropy apply to systems in non-equilibrium?
@@
At the outset, we stress that the above issue arises only when the system is not in thermal equilibrium. In the thermal equilibrium, as discussed earlier in Sec. III.1, the VN entropy and the TH entropy are equivalent (under certain general conditions). Before we discuss two different approaches to resolve the problem, we first note that the VN entropy is the quantum-analog of the Gibbs (GB) entropy [17]. Hence, some of the major criticisms faced by the GB entropy are also applicable to the VN entropy. For example, the GB entropy of a system is time invariant since the “shape” of the probability distribution function does not change under Hamiltonian evolution of the system [7]. There are two different ways one can tackle this time-invariance problem. First, one considers that the VN entropy formula needs modification. Accordingly, new entropy formula is proposed which increases as the system evolves with time (this entropy increases and finally saturates in equilibrium). The observational entropy is one such popular alternative to the VN entropy [14, 18]. Second, instead of considering that the VN entropy formula is wrong, one may consider that, while the formula is correct, the time evaluation of density matrix $\rho$ is not correctly described by von Neumann-Liouville equation. We prefer the second option for the following reasons.
@@
Like the TH entropy, the VN entropy is a state function - it depends only on $\rho$ and does not depend on any choice or preference of an observer. The VN entropy formula correctly captures equilibrium thermodynamics of a system. In fact, starting with the VN entropy formula, one can derive the expression of the equilibrium Gibbs distribution using the maximum entropy principle and some other conditions [19]. All these point out that the expression of the VN entropy is a suitable one to study the thermodynamics of a quantum system; main challenge here is to figure out how one the the density matrix evolves from a given initial state to the final one (representing thermal equilibrium). In the present study, we leave out any such detail discussion on how the density matrix evolves.
@@
The second major objection against the VN entropy is that it obeys the subadditivity property while the TH entropy is additive (for a large uniform system where the interaction energy between its two parts can be ignored). The subadditivity property is stated as follows: if a system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices of the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [20]. This implies that, if the system is in a pure state, its VN entropy will be zero but the sum of the VN entropies of its parts is positive (the sum can be zero in a very special case). This contradicts the notion of additivity of the TH entropy. We can circumvent this apparent contradiction in the following way.
As we noted earlier, the VN entropy is the quantum-analog of the GB entropy. The GB entropy is applicable for an open system which is in thermal equilibrium with its bath. For such a thermodynamic system, its TH entropy is equal to the TH entropies of its parts. Now to verify if the VN entropy has this additivity property, we first note that, the quantum thermalization is defined with respect to a smaller subsystem of a large system [21, 22, 1, 5, 23]. Here, a part of the system is thermalized while the rest of the system works as its bath. Therefore, instead of studying additivity property of the whole quantum system, we should study it for a subsystem. It is know that the VN entropy follows a volume law for excited states (especially the states from the middle of its spectrum) [24, 25, 26, 27]. In fact, for our system (see Sec. IV.3), we see that the average VN entropy $\overline{S}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size, as seen in Fig. 1.
Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}_{VN}$) grows linearly with subsystem size. Here total system size is $N=16$.
@@ IV Entropy of large quantum systems
We already saw in Sec. II and Sec. III.1 that for a thermalized large isolated or open quantum system, the VN entropy is equivalent to the TH entropy (or to an appropriate quantum statistical entropy). It is non-trivial to check this hypothesis for a small subsystem of a large quantum system. In this section we check, mainly through numerical analysis, whether and how the VN entropy is equivalent to the TH entropy for a subsystem. For completeness and more clarity, we first briefly summarize how the VN entropy is equal (or equivalent) to the Th entropy for a thermalized isolated and open system. We exclude from our discussion the systems which are not in equilibrium or the systems which do not thermalize.
Figure 2: The subsystem VN entropy ($S_{VN}$) for individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system. Here total system size is 16 and subsystem size is 6.
@@ IV.1 Isolated quantum system
As discussed in Sec. II.3, an isolated quantum system will be in a mixed state, for one or multiple reasons. Using the principle of equal a priori probabilities, one can show that the VN entropy is equal to the qBN entropy (which in turn is equivalent to the TH entropy).
Assume that we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5]. Let {|$\{\ket{n}\}$ be the set of the eigenkets of the system Hamiltonian, and $E_{n}$ be the eigenvalue corresponding to |$\ket{n}$. The system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E,E+\Delta E]$. Accordingly, the thermodynamic entropy ($S_{TH}$), or equivalently, the quantum Boltmann entropy ($S_{qBN}$) is,
@@
Now to calculate the corresponding von Neumann entropy ($S_{VN}$) of the isolated quantum system, we note that the system is in a mixed state, represented by the following density matrix,
$\rho=\sum_{\begin{subarray}{c}n\\ E_{n}\in\left(E,E+\Delta E\right]\end{subarray}}p_{n}\ket{n}\bra{n},$ (9)
where $p_{n}$ is the probability to find the system in state |$\ket{n}$. Since $\Delta E$ is macroscopically small, all the accessible microstates (here, eigenkets) will essentially have almost the same energy. Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{D}$. As $S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)$, we need the eigenvalues of $\rho$ to efficiently compute $S_{VN}$. It is clear from Eq. 9 that |$\ket{n}$ is an eigenket of $\rho$ with eigenvalue $\frac{1}{D}$, i.e., $\rho\ket{n}=\frac{1}{D}\ket{n}$ for any energy eigenket from the shell. This implies that
$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho)=k_{B}\ln D.$ (10)
We see from Eqs. 8 and 10 that, for an isolated quantum system, the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$), provided that the system is large and in thermal equilibrium. The detail discussion on the conditions under which this equivalence works can be found in Sec. II.2.

### v2

@@ III.1 As basis-independent representation of quantum statistical entropy
One way to understand the VN entropy is to view it as the basis-independent representation of the quantum statistical entropy of a thermodynamic system in equilibrium. The VN entropy for an appropriate density matrix can give us the qBN or qGB entropy. In fact, while deriving the expression of the thermodynamic (TH) entropy of a subsystem, Landau also came up with the same expression that appears in Eq. 3 as the basis-independent representation of the quantum statistical entropy [9].
For an isolated system, described by a microcanonical (mc) ensemble, we get the following from Eq. 1: $S_{qBN}=-k_{B}\ln D=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})=S_{VN}$, where $\rho_{mc}=\frac{1}{D}I$ with $D$ being the dimension of the relevant subspace of the Hilbert space of the system and $I$ is the identity matrix. So we find here that the VN entropy can be viewed as a basis-independent representation of the qBN entropy of an isolated system in thermal equilibrium.
Similarly, for an open system, we get the following from the Eq. 2, $S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}=-k_{B}\text{Tr}(\rho_{\beta}\ln\rho_{\beta})=S_{VN}$, where $\rho_{\beta}=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$. Here $H$ and $\beta$ are, respectively, the Hamiltonian and the inverse temperature of the open system. We note again that the VN entropy of an open system in equilibrium can be viewed as a basis-independent representation of the qGB entropy.
@@ III.2 As entropy of thermodynamic ensemble
Sometimes we may not know in which pure state our quantum system is in; von Neumann introduced the concept of the density matrix ($\rho$) to describe the physics of such system [8]. The density matrix is a statistical mixture of many pure states; for example,
@@
While following the Gibbsian way, one immediately faces problems in defining entropy of the system in the state $\rho$. First of all, the decomposition in Eq. 4 is not unique and, then, the pure states {|$\{\ket{\psi_{i}}\}$ in a decomposition may not be in general distinguishable (as they are not in general orthonormal). Through thermodynamic considerations, von Neumann proposed a formula (as given in Eq. 3) to quantify the entropy of a system in state $\rho$. We will not go through his arguments here; rather, we will discuss two alternative ways to understand the VN entropy. Before that, we briefly discuss below the two main quantum processes and their effects on the VN entropy.
As discussed by von Neumann, we can consider two basic processes that a system may undergo: (1) time evaluation of the system, and (2) measurement on the system. If $\rho(0)$ is the density matrix at time $t=0$ then the density matrix at the $t$ is given by $\rho(t)=U(t)\rho(0)U^{{\dagger}}(t)$, where $U(t)$ is the time evaluation operator. It is not difficult to see that the VN entropy is invariant under this time evaluation process, i.e., $S_{VN}(t)=S_{VN}(0)$. von Neumann showed that a measurement, in general, irreversibly changes the density matrix to another density matrix $\rho\to\rho^{\prime}$, where
@@
We note that the set of states {|$\{\ket{\phi_{n}}\}$ are the eigenkets of the observable for which one is performing the measurement and the projection operator $P_{n}=\ket{n}\bra{n}$ |. The subtleties in dealing with degeneracy can be found in von Neumann’s work [8]. He showed that the entropy of a system, in general, increases (or remains constant but never decreases) due to the measurement process, i.e., $S_{VN}(\rho^{\prime})\geq S_{VN}(\rho)$. If we identify the first process (time evaluation) and the second process (measurement) with the thermodynamic reversible and irreversible processes, respectively, then these results are consistent with the thermodynamic laws. In the following, we now discuss two different ways one can intuitively understand the entropy formula.
As in Sec. II.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero (see Eq. 1). In this case, as we see from Eq. 3, the VN entropy is also zero. When the system is not in pure state, rather in a statistical mixture of pure states (represented by $\rho$), a formula for entropy must quantify the amount of “mixedness” (or statistical impurity) in the state $\rho$, similar to how the BN entropy quantifies the amount of “uncertainty” of a system to be found in a microstate. If we now see the decomposition in Eq. 4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-k_{B}\sum_{i}p_{i}\ln p_{i}$, which is the Shannon (SN) entropy for the probability distribution $\{p_{i}\}$. Unfortunately, the decomposition of a mixed state is not unique, hence this way of defining entropy is unacceptable. One way to circumvent this problem is to consider all possible decompositions of $\rho$ and take the one for which $S_{SN}(\{p_{i}\})$ is minimum. The advantage of defining entropy this way is that the entropy is now unique for a given $\rho$. In fact it can be proved that the VN entropy of a given $\rho$ corresponds to the aforementioned minimum [17, 18]:
@@
Another way to understand the VN entropy is as follows. If we take a set of orthonormal states to be a set of distinguishable microstates, we can define an entropy from the probabilities associated with these microstates. If {|$\{\ket{\phi_{n}}\}$ is a set of orthonormal states, the entropy one can define is $S_{SN}(\{w_{i}\})=-k_{B}\sum_{n}w_{n}\ln w_{n}$, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. Unfortunately, again, this entropy is not unique as it depends on the set of orthonormal states (microstates). To understand how to avoid this problem, we first define a new density matrix $\rho^{\prime}=\sum_{n}w_{n}\ket{\phi_{n}}\bra{\phi_{n}}$ |, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. We may note that |$\ket{\phi_{n}}$ is an eigenket of $\rho^{\prime}$ with eigenvalue $w_{n}$. This density matrix ($\rho^{\prime}$) is in the same form as the one defined in Eq. 5. For this updated density matrix after measurement ($\rho\to\rho^{\prime}$), as mentioned earlier, von Neumann showed that $-\text{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\text{Tr}(\rho\ln\rho)$ [8]. Thus we see that the VN entropy corresponds to the minimum of $S_{SN}(\{w_{i}\})$, which we obtain when we choose the microstates as the eigenkets of the given density matrix $\rho$. Accordingly,
@@
III.3 Addressing objections against VN entropy
@@
Time-invariance of VN entropy:
@@
The first of these objections is that the VN entropy of an isolated quantum system does not change (more specifically, increase) as the system evolves (unitarily) with time. As seen earlier, $S_{VN}(t)=S_{VN}(0)$ when $\rho(0)\to\rho(t)$ under the time evaluation operator $U(t)$. It is known that, for an isolated system, the TH entropy increases as the system approaches thermal equilibrium. Does not the VN entropy apply to systems in non-equilibrium?
At the outset, we stress that the above issue arises only when the system is in a non-equilibrium state and we want to know the entropy during a process. In the thermal equilibrium, as discussed earlier in Sec. III.1, the VN entropy and the TH entropy are equivalent (under certain general conditions). So, the VN entropy is capable of determining the change in (TH) entropy due to a process (the system is in equilibrium before and after the process).
We also note that this time-invariance of VN entropy is similar to that of the GB entropy. The GB entropy of a system is time-invariant since the “shape” of the probability distribution function does not change under the Hamiltonian evolution of the system [7, 19]. It may be noted that the similarity of the problems faced by the VN and GB entropies is not unexpected, since the VN entropy is the quantum analog of the Gibbs (GB) entropy [20]. There are many attempts in resolving this issue of time-invariance of the GB entropy under the Hamiltonian evaluation; discussions on them may be found elsewhere [7, 19].
@@
While trying to resolve the time-invariance problem of the VN entropy, one can think of two major different approaches. First, one considers that the VN entropy formula needs modification (or some kind of generalization) so that the entropy from the new formula increases (and saturates in equilibrium) as the system evolves with time. The observational entropy is one such popular alternative to the VN entropy [16, 21, 22]. Second, instead of considering that the VN entropy formula is wrong, one may consider that the time evaluation of the density matrix $\rho$ is not correctly described by the von Neumann-Liouville equation. We prefer the second option for the following reasons.
@@
Like the TH entropy, the VN entropy is a state function in the sense that it depends only on the state $\rho$ of the system and does not depend on any choice or preference of an observer. The VN entropy formula correctly captures the equilibrium thermodynamics of a system (as mentioned, in equilibrium, $S_{VN}\equiv S_{TH}$). In fact, starting with the VN entropy formula, one can derive the expression of the equilibrium Gibbs distribution using the maximum entropy principle and some other conditions [23]. All these point out that the expression of the VN entropy is a suitable one to study the thermodynamics of a quantum system; main challenge here is to figure out how the density matrix evolves from a given initial state to the final one (representing thermal equilibrium). In the present study, we leave out any such detailed discussion on how the density matrix evolves in an irreversible process. We just mention here in passing that, although finding an evaluation equation of $\rho$ for an isolated system undergoing an irreversible process can be challenging, such evaluation for an open system is well studied via the Gorini-Kossakowski-Sudarshan-Lindblad equations [24]. The last thing we may want to note here is that, in some special cases, two approaches mentioned above can actually be equivalent. For example, for single coarse-graining, the observational entropy is actually equal to the VN entropy of the coarse-grained density matrix [21, 25].
@@
The second major objection against the VN entropy is that it obeys the subadditivity property while the TH entropy is additive (for a large uniform system where the interaction energy between its two parts can be ignored). The subadditivity property is stated as follows: if a bipartite system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices of the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [26]. This implies that, if the system is in a pure state, its VN entropy will be zero but the sum of the VN entropies of its parts is positive (the sum can be zero in a very special case). This contradicts the notion of the additivity property that the TH entropy has. We can circumvent this apparent contradiction in the following way.
To verify whether the VN entropy has the additivity property, we first note that the quantum thermalization is defined with respect to a subsystem of a (much) larger system [27, 28, 1, 5, 29, 2]. Essentially one here studies whether the subsystem is thermalized while the rest of the system works as its bath. Therefore, instead of studying the additivity property of the whole quantum system, we should study it for a subsystem. This is exactly what one does for the classical case where one studies additivity property of the GB entropy (TH entropy) of an open system in thermal equilibrium with its bath. For an open system, the additivity implies that the entropies of its parts sum up to the entropy of the system itself (the entropy of the bath does not play any role here).
With the above understanding, we now check if the VN entropy of a subsystem has the additivity property. It is known that the VN entropy follows a volume law for excited states (especially the states from the middle of its spectrum) [30, 31, 32, 33]. In fact, for our system with $\Delta_{2}=0.5$ (see Sec. IV.3), we see that the average VN entropy $\overline{S}^{sb}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size; relevant results can be found in Fig. 1.
Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}^{sb}_{VN}$) grows linearly with subsystem size. Here total system size is $N=16$ and $\Delta_{2}=0.5$.
@@ IV Entropy of large quantum systems
In Sec. III.1, we have seen that the VN entropy of a thermalized isolated or open system is a basis-independent representation of the quantum statistical entropy (which, under certain conditions, is equivalent to the TH entropy). There we start with the qBN or qGB entropy and then show that it is simply equal to the VN entropy corresponding to some density matrix (we were not much concerned about any significance or possible interpretation of the density matrix).
In this section, we go in the opposite direction: we first calculate the VN entropy independently and then verify if that is equal (equivalent) to the TH entropy of the system (subsystem). To calculate the VN entropy, we take an appropriate density matrix which describes the system (subsystem) in thermodynamic equilibrium.
In the following, we discuss the issue of equivalence separately for the isolated, open and subsystem (of a much larger system).
@@ IV.1 Isolated quantum system
We consider an isolated quantum system in thermal equilibrium. As discussed in Sec. II.3, a large isolated quantum system will be in a mixed state, for one or multiple reasons. An isolated system in thermal equilibrium is appropriately described by a microcanonical (mc) ensemble. Assume that we prepare the system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5]. Let {|$\{\ket{n}\}$ be the set of eigenkets of the Hamiltonian of the system, and $E_{n}$ be the eigenvalue corresponding to |$\ket{n}$. The density matrix which describes the isolated system is given by,
$\rho_{mc}=\sum_{\begin{subarray}{c}n\\ E_{n}\in\left(E,E+\Delta E\right]\end{subarray}}p_{n}\ket{n}\bra{n},$ (8)
where $p_{n}$ is the probability to find the system in state |$\ket{n}$. Since $\Delta E$ is macroscopically small, all the accessible microstates (here, the eigenkets) will essentially have almost the same energy. Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{D}$ (see aslo Sec. II.3). As $S_{VN}=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})$, we need the eigenvalues of $\rho_{mc}$ to efficiently compute $S_{VN}$. It is clear from Eq. 8 that |$\ket{n}$ is an eigenket of $\rho_{mc}$ with eigenvalue $\frac{1}{D}$, i.e., $\rho_{mc}\ket{n}=\frac{1}{D}\ket{n}$ for any energy eigenket from the shell. This implies that
$S_{VN}=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})=k_{B}\ln D.$ (9)
If we now directly calculate the TH entropy (or equivalently, the qBN entropy), following our discussion in Sec. II.3, it would be
@@
We see from Eqs. 9 and 10 that, for an isolated quantum system, the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$), provided that the system is large and in thermal equilibrium. The detailed discussion on the conditions under which this equivalence works can be found in Sec. II.2.